# PPAD: Iterative Interactions of Prediction and Planning for End-to-end Autonomous Driving

### v3

@@ Abstract
We present a new interaction mechanism of prediction and planning for end-to-end autonomous driving, called PPAD (Iterative Interaction of Prediction and Planning Autonomous Driving), which considers the timestep-wise interaction to better integrate prediction and planning. An ego vehicle performs motion planning at each timestep based on the trajectory prediction of surrounding agents (e.g., vehicles and pedestrians) and its local road conditions. Unlike existing end-to-end autonomous driving frameworks, PPAD models the interactions among ego, agents, and the dynamic environment in an autoregressive manner by interleaving the Prediction and Planning processes at every timestep, instead of a single sequential process of prediction followed by planning. Specifically, we design ego-to-agent, ego-to-map, and ego-to-BEV interaction mechanisms with hierarchical dynamic key objects attention to better model the interactions. The experiments on the nuScenes benchmark show that our approach outperforms state-of-the-art methods.
@@ 2.2 End-To-End Autonomous Driving
Recently, more and more works have focused on end-to-end autonomous driving due to its merits in reducing internal accumulative errors and direct yet simple learning objectives. Typical methods [44, 2] take the raw visual inputs to regress the final control command or trajectory points without any view transformations. To embrace reinforcement learning, a series of following works [53, 44, 8] utilize policy-based or valued-based approaches to improve driving behavior. With the popularity of BEV representation, more advanced architectures [21, 9] are introduced to attach more interpretability and help deal with the complex interactions in the driving scenarios. Another merit of BEV representation lies in its simplicity in fusing multi-modality sensors. Moreover, the modularized approaches [18, 22, 17] decouple the end-to-end learning-based methods into several submodules or subtasks, while in a multi-task learning manner. The unified design could propagate and share learning context between modules through queries or feature maps.
@@ 3.2.2 Planning Process
$\mathbf{E}^{k}=\sum_{s\in\mathbf{S}}{\texttt{MHCA}(\mathbf{E},\mathbf{A}^{k},\mathcal{M}(p^{t}_{\mathbf{E}},p^{t+1}_{\mathbf{A}^{k}},s))},\small{k\in[1,N_{A}^{mot}]},$ (1)
@@ 3.5 End-to-End Learning
4 Experiments
4.1 Experimental Setup
We conduct experiments on the publicly available nuScenes dataset [3], which is one of the largest and most popular datasets designed for autonomous driving tasks, including 2D/3D object detection, tracking, segmentation, planning, and prediction. We also evaluate the proposed components to verify their effectiveness through ablation studies.
Dataset The nuScenes dataset [3] provides about $1$K $20$-second diverse driving scenes collected in Boston, Pittsburgh, Las Vegas, and Singapore for open-loop settings. Key samples which contain $6$ camera images are annotated at $2$Hz.
The Argoverse2 dataset [43] comprises $1$K scenes, each spanning $15$-second, providing 3D vectorized maps and $10$Hz 3D cuboid annotations for objects. It captures $7$ RGB images on each frame in comparison to $6$ images in nuScenes [3]. We align the data sampling frequency to nuScenes [3] by downsampling the annotated frames by an interval of $5$ frames.
Metrics We adopt the widely used metrics Displacement Error and Collision Rate as evaluation criteria [17]. Displacement error is measured with ground-truth trajectories over the time steps $1$-s, $2$-s, $3$-s. Collision Rate (CR) measures how often the collision occurs between the ego vehicle and the other agents along the planning horizon. We notice that the UniAD [18] adheres to different calculations in comparison to ST-P3 [17] and VAD [22]: the former [18] reports the evaluations at each second. In contrast, the latters [17, 22] reports the results of the cumulative average by each second. We faithfully make comparisons of our method to others with these two kinds of computations.
Implementation Details We follow the standard settings as proposed by UniAD [18] and VAD [22]. The perception range for the $x$-axis is [-15m, 15m] and [-30m, 30m] for the $y$-axis. Note that the planning trajectories for the ego and the agents have no limitation in range. We adopt ResNet-50 [14] as the image backbone and BEVFormer [28] as the view transformation layer to aggregate multi-camera features. The perception transformer recurrently encodes $2$-s history information. Then, our PPAD is targeted to plans for the future $3$-s trajectory for the ego vehicles. It also conducts the tasks of motion prediction and map construction. Our PPAD trains with a batch size of $1$ using the Adam [24] optimizer set with an initial learning rate of 2e-4.
4.2 Main Results
Planning Results We faithfully compare our PPAD with current state-of-the-art methods in nuScenes [3] dataset using the metrics from ST-P3 [17] and the metrics from UniAD [22]. As shown in Tab. 1, our PPAD outperforms the current state-of-the-art performance by a large margin. Especially for the L2 distance metrics, there are about $20$% of consistent improvements can be observed along the temporal horizon. Thanks to the iterative interaction of prediction and planning, PPAD can help avoid collisions, leading to better results on collision rate compared to the one-shot representative VAD [22]. At the same time, we maintain a competitive efficiency compared to UniAD [18].
We further make a fair comparison between our method and the baseline, VAD [22], on the other dataset of Argoverse2 [6, 43]. It is observed that we consistently achieve superior results on both displacement error and collation rate metrics. As shown in Tab. 2, our method can consistently outperform the baseline with a clear margin in both L2 distance and collision rate metrics.
@@ 4.2 Main Results
Table 1: Open-loop planning results on the nuScenes dataset [3]. The results of other methods are obtained from the original paper. We faithfully re-evaluate VAD [22] based on the UniAD [18] metrics. As for our PPAD, we provide two versions of results that utilize different training strategies. Progress. means that we follow the progressive training pipeline as proposed in VAD [22], which trained all tasks except the planning task in the first 48 epochs and then finetuned with another 12 epochs for the planning task. The second row for our method trains the whole network for 60 epochs and then finetuned for another 12 epochs incorporating noisy trajectories.
@@
Table 2: Open-loop planning results on the Argoverse2 dataset [3]. The results are evaluated under the ST-P3 [17] and VAD [22] metrics.
@@ 4.3 Ablation Study
The following experiments adhere to the progressive training pipeline as proposed in VAD [22] on the nuScenes dataset [3].
Effectiveness of Designs We provide ablation studies to verify the effectiveness of our proposed components. As shown in Tab. 4, the proposed PPAD framework (row 2) brings a remarkable and significant improvement compared with one-shoot methods [22] (row 1). The multi-step interactions help the ego agent better understand the intention and potential effects brought by its actions along the temporal horizon, leading to an over $10$% L2 distance error reduction. We can observe a further improvement regarding the L2 distance with our proposed key objects attention (row 3). The slight degradation in collision rate might be due to the key objects attention being conducted on the ego with each mode of the agents, while the diverse modes of the agents mislead the behavior of the ego vehicle. From rows 4-6, the confidence-aware collision loss and noisy trajectory as the prediction are proposed for circumventing this phenomenon and further escalate the capability in both planning accuracy and the ability to avoid a collision.
  PPAD Key Objects Attn. $\mathcal{L}_{CA-Col}$ Noisy Traj. L2 (m) $\downarrow$ Collision (%) $\downarrow$
@@
Effectiveness of Interactions Our PPAD framework enables richer interactions among scene contexts, introducing local and global understandings of the world to the model. We further conduct ablation studies to demonstrate the performance gains brought by the interactions. We conduct the ablation study on interactions under the setting (Tab. 4 in row 4) without using the noisy trajectory as the prediction for better comparison. As illustrated in Tab. 5, we can achieve the best performance in terms of L2 distance and collision by incorporating all of the interactions.
@@
  Interaction Iterations L2 (m) $\downarrow$ Collision (%) $\downarrow$
@@
   2 (Every 1.5 sec) 0.36 0.63 0.96 0.65 0.14 0.20 0.43 0.25
3 (Every 1.0 sec) 0.37 0.63 0.95 0.65 0.10 0.15 0.39 0.21
   6 (Every 0.5 sec) 0.31 0.56 0.87 0.58 0.08 0.12 0.38 0.19
@@
Effect of Key Objects Attention Distance Ranges The hierarchical dynamic key objects attention is proposed to understand the the dynamic surrounding driving agents and the environments in a coarse-to-fine manner. Tab. 7 shows the ablation study, which was conducted by setting different distance ranges. We use the distance range set of $\{+\infty\ m,15\ m,7.5\ m\}$, which achieves the best performance in Tab. 7, as our model’s default setting.
  Distance Range L2 (m) $\downarrow$ Collision (%) $\downarrow$
1s 2s 3s Avg. 1s 2s 3s Avg.
  $\{+\infty\ m,7.5\ m,3.75\ m\}$ 0.33 0.56 0.87 0.59 0.08 0.12 0.35 0.19
$\{+\infty\ m,15\ m,7.5\ m,3.75\ m\}$ 0.34 0.59 0.91 0.61 0.11 0.18 0.41 0.23
  $\{+\infty\ m,15\ m,7.5\ m\}$ 0.31 0.56 0.87 0.58 0.08 0.12 0.38 0.19
Table 7: The ablation study of different distance ranges for conducting the hierarchical dynamic key objects attention.

### v4

@@ Abstract
We present a new interaction mechanism of prediction and planning for end-to-end autonomous driving, called PPAD (Iterative Interaction of Prediction and Planning Autonomous Driving), which considers the timestep-wise interaction to better integrate prediction and planning. An ego vehicle performs motion planning at each timestep based on the trajectory prediction of surrounding agents (e.g., vehicles and pedestrians) and its local road conditions. Unlike existing end-to-end autonomous driving frameworks, PPAD models the interactions among ego, agents, and the dynamic environment in an autoregressive manner by interleaving the Prediction and Planning processes at every timestep, instead of a single sequential process of prediction followed by planning. Specifically, we design ego-to-agent, ego-to-map, and ego-to-BEV interaction mechanisms with hierarchical dynamic key objects attention to better model the interactions. The experiments on the nuScenes benchmark show that our approach outperforms state-of-the-art methods. Project page at https://github.com/zlichen/PPAD.
@@ 2.2 End-To-End Autonomous Driving
Recently, more and more works have focused on end-to-end autonomous driving due to its merits in reducing internal accumulative errors and direct yet simple learning objectives. Typical methods [45, 2] take the raw visual inputs to regress the final control command or trajectory points without any view transformations. To embrace reinforcement learning, a series of following works [54, 45, 8] utilize policy-based or valued-based approaches to improve driving behavior. With the popularity of BEV representation, more advanced architectures [22, 9] are introduced to attach more interpretability and help deal with the complex interactions in the driving scenarios. Another merit of BEV representation lies in its simplicity in fusing multi-modality sensors. Moreover, the modularized approaches [19, 23, 18, 14] decouple the end-to-end learning-based methods into several submodules or subtasks, while in a multi-task learning manner. The unified design could propagate and share learning context between modules through queries or feature maps.
@@ 3.2.2 Planning Process
$\mathbf{E}^{k}=\sum_{s\in\mathbf{S}}{\texttt{MHCA}(\mathbf{E},\mathbf{A}^{k},\mathcal{M}(p^{t}_{\mathbf{E}},p^{t+1}_{\mathbf{A}^{k}},s))},k\in[1,N_{A}^{mot}],$ (1)
@@ 3.5 End-to-End Learning
Table 1: Open-loop planning results on the nuScenes dataset [3]. The results of other methods are obtained from the original paper. We faithfully re-evaluate VAD [23] based on the UniAD [19] metrics. As for our PPAD, we provide two versions of results that utilize different training strategies. Progress. means that we follow the progressive training pipeline as proposed in VAD [23], which trained all tasks except the planning task in the first 48 epochs and then finetuned with another 12 epochs for the planning task. The second row for our method trains the whole network for 60 epochs and then finetuned for another 12 epochs incorporating noisy trajectories. The latency of ST-P3 [18], VAD [23], and ours are measured on one NVIDIA Geforce RTX 3090 GPU, while UniAD is measured on one NVIDIA Tesla A100 GPU.
@@
Table 2: Open-loop planning results on the Argoverse dataset [3]. The results are evaluated under the ST-P3 [18] and VAD [23] metrics.
@@
4 Experiments
4.1 Experimental Setup
Dataset We evaluate our method on two challenging large-scale real-world datasets, nuScenes [3] and Argoverse2 [44]. We conduct ablation studies to evaluate the effectiveness of our proposed components on the nuScenes [3] dataset. The nuScenes dataset [3] provides about $1$K $20$-second diverse driving scenes collected in Boston, Pittsburgh, Las Vegas, and Singapore for open-loop settings. Key samples which contain $6$ camera images are annotated at $2$Hz. The Argoverse2 dataset comprises $1$K $15$-second scenes. It captures $7$ RGB images on each frame. We align the data sampling frequency to nuScenes [3] by downsampling the annotated frames by an interval of $5$ frames.
Metrics We adopt the metrics of L2 Displacement Error (L2) and Collision Rate (CR) for evaluation [18]. L2 is measured between the prediction and ground-truth trajectories over the timesteps $1$-s, $2$-s, and $3$-s, evaluating the trajectory quality. Collision Rate (CR) measures how often the collision occurs between the ego vehicle and the other agents along the planning horizon, reflecting the trajectory safety. We notice that the UniAD [19] adheres to different calculations from ST-P3 [18] and VAD [23]: the former [19] reports the evaluations at each second. In contrast, the latters [18, 23] reports the results of the cumulative average by each second. We faithfully make comparisons of our method to others with these two kinds of computations.
Implementation Details We strictly follow the standard settings as proposed by UniAD [19] and VAD [23], which did not use the information of the historical ego trajectory. Same as VAD [23], the perception range is $60m\times 30m$. PPAD also recurrently encodes $2$-s historical information into BEV and predicts the $3$-s trajectory in the future. It also conducts the tasks of motion prediction and map construction. Our PPAD trains with a batch size of $1$ using the Adam [25] optimizer set with an initial learning rate of 2e-4. It takes us about six days to conduct the end-to-end training on eight NVIDIA A30 24GB GPUs.
4.2 Main Results
Planning Results As shown in Tab. 1, our PPAD outperforms the current state-of-the-art performance by a large margin. Especially for the L2 distance metrics, there are about $20$% of consistent improvements can be observed along the temporal horizon. Thanks to the iterative interaction of prediction and planning, PPAD can help avoid collisions, leading to better results on collision rate compared to the one-shot representative VAD [23]. At the same time, we maintain a competitive efficiency compared to UniAD [19].
We further make a fair comparison between our method and the baseline, VAD [23], on the other dataset of Argoverse 2 [6, 44]. In Tab. 2, our method can consistently outperform the baseline with a clear margin in both L2 distance and collision rate metrics.
@@ 4.3 Ablation Study
The following experiments adhere to the progressive training pipeline as proposed in VAD [23].
Effectiveness of Designs We provide ablation studies to verify the effectiveness of our proposed components. As shown in Tab. 4, the proposed PPAD framework (row 2) brings a remarkable improvement compared with one-shoot methods [23] (row 1). The multi-step interactions help the ego agent better understand the intention and potential effects brought by its actions along the temporal horizon, leading to an over $10$% L2 distance error reduction. We can observe a further improvement regarding the L2 distance with our proposed key objects attention (row 3). The slight degradation in collision rate might be due to the key objects attention being conducted on the ego with each mode of the agents, and the diverse modes of the agents mislead the behavior of the ego vehicle. From rows 5-7, the confidence-aware collision loss and noisy trajectory as the prediction can circumvent this phenomenon and further escalate the capability in planning accuracy and avoiding a collision. When we keep all of the remaining components while not applying key objects attention (row 4), we observed the degradations from row 7 in L2 (0.58 m vs. 0.59 m) and collision rate (0.19% vs. 0.21%), proving the effectiveness of the key objects attention.
  PPAD Key Objects Attn. $\mathcal{L}_{CA-Col}$ Noisy Traj. L2 (m) $\downarrow$ Collision (%) $\downarrow$ Latency (ms)
@@
Effectiveness of Interactions Our PPAD framework enables richer interactions among scene contexts, introducing local and global understandings of the world to the model. We further conduct ablation studies to demonstrate the performance gains brought by the interactions. We conduct the ablation study on interactions under the setting (Tab. 4 in row 5) without using the noisy trajectory as the prediction for better comparison. As illustrated in Tab. 5, we can achieve the best performance in terms of L2 distance and collision by incorporating all of the interactions.
@@
  Interaction Iterations L2 (m) $\downarrow$ Collision (%) $\downarrow$ Latency (ms)
@@
   2 (Every 1.5 sec) 0.36 0.63 0.96 0.65 0.14 0.20 0.43 0.25 306
3 (Every 1.0 sec) 0.37 0.63 0.95 0.65 0.10 0.15 0.39 0.21 326
   6 (Every 0.5 sec) 0.31 0.56 0.87 0.58 0.08 0.12 0.38 0.19 385
@@ 5 Conclusion
Acknowledgements
The authors are thankful for the financial support from the Hetao Shenzhen-HongKong Science and Technology Innovation Cooperation Zone (HZQB-KCZYZ-2021055), this work was also supported by Shenzhen Deeproute.ai Co., Ltd (HZQB-KCZYZ-2021055).